OpenStudy (anonymous):

The area of a rectangle is represented by x^2-2x-8 find an expression for the length of each side if x=5

OpenStudy (some.random.cool.kid):

hmmm

myininaya (myininaya):

sounds like you need to factor x^2-2x-8 as a first step

myininaya (myininaya):

area=x^2-2x-8 LW=( )( ) <--this is the factored form of the area

myininaya (myininaya):

so what two numbers multiply to be -8 and add up to be -2?

myininaya (myininaya):

@Ednac are you there?

OpenStudy (anonymous):

there is nothing to factor from the initial problem @myininaya

myininaya (myininaya):

yes you need to factor x^2-2x-8

myininaya (myininaya):

to find the length and the width of the rectangle

OpenStudy (anonymous):

it is as factored as it gets

myininaya (myininaya):

-4(2)=-8 -4+2=-2 it can be factored

myininaya (myininaya):

so given that the factored form of x^2-2x-8 is ?

OpenStudy (anonymous):

is there a term that can be taken out fropm all of the terms other than one?

OpenStudy (anonymous):

Where are you going with the fact that it is factorable

myininaya (myininaya):

to find the length and width of the rectangle with area=x^2-2x-8

myininaya (myininaya):

Area of rectangle=Length * Wdith x^2-2x-8 =( )* ( )

myininaya (myininaya):

now I basically told you the factored form

myininaya (myininaya):

since I said -4(2)=-8 and -4+2=-2

myininaya (myininaya):

so x^2-2x-8=(x-4)(x+2)

myininaya (myininaya):

now you have the length and the width

myininaya (myininaya):

your question wants to know the length and the width when x=5

OpenStudy (anonymous):

awesome job at explaining @myininaya

myininaya (myininaya):

I tried (or try)